b) Read Section 2.3.6 (on page 161) of the 8'th edition of Rosen's book on partial functions. Let A and B be finite sets, with |A| = m and |B| = n. Calculate the number of partial functions f: A -> B.
1
Expert's answer
2020-12-18T14:49:51-0500

A partial function f rom "A" to "B" is a map from "X \subset A" to "B" . If "X" has "k" elements such that "0 \leq k \leq n" . So there are "n^k" of such map.

In addition, there will be "\begin{pmatrix}
 m\\
 k
\end{pmatrix}" subsets of "A". So, the number of partial functions will be;


"\sum_{k=0}^m\begin{pmatrix}
 m\\
 k
\end{pmatrix}n^k=(1+n)^m"